# Diffusion approximation of the stationary distribution of a two-level single server queue

### v3

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study in the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation. It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [1, 2, 11, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
@@
Note: This version corrects Section 5.1 in addition to fixing typos.
11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan & School of Data Science, Chinese University of Hong Kong, Shenzhen, China.
@@ 1 Introduction
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. It may be considered to approximate using a quasi-birth-and-death process, and it is not hard to numerically compute its performance. However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential. Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of this 2-level $GI/G/1$ queues under a certain scaling in heavy traffic. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@ 2 Framework for the 2-level GI/G/1 queues
We take the stochastic framework introduced in [12] to study the 2-level $GI/G/1$ queue. Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining service time, respectively, at time $t$. Define
@@ 3.1 Assumptions
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov. Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with respect to the time-shift operator $\Psi$. We assume the following assumptions.
@@
Furthermore, (3.b) and (3.c) imply that $1-\rho^{(r)}_{i}=rb_{i}/c_{i}+o(r)$, $\rho_{i}\equiv\frac{\lambda_{i}}{c_{i}\mu}=1$ and $\lambda_{i}=c_{i}\mu$. Hence, $\lambda_{1}=\lambda_{2}$ if and only if $c_{1}=c_{2}$. It is easy to see that $X^{(r)}(t)$ has the stationary distribution by (3.e). This distribution is denoted by $\nu^{(r)}$.
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the stability condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
@@ 3.2 Main results
    $\lim_{r\downarrow 0}\nu_{i|cd}^{(r)}(B)=\widetilde{\nu}_{icd}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.6)
@@
$\rho^{(r)}=\alpha^{(r)}_{e}\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}S^{(r)}_{k}.$ (3.12)
@@
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.13)
@@
Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{M}$ which is defined as $\widetilde{\nu}_{M}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{1}(M)+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{2}(M)$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
$\widetilde{A}_{1}(M)=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
and $\widetilde{A}_{2}(M)=1-\widetilde{A}_{1}(M)$.
@@
Since $\sigma_{i}^{2}=2$ and $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$ in the exponential case, $\widetilde{\nu}_{M}$ and $\widetilde{A}_{i}(M)$ agree with $\widetilde{\nu}$ and $\widetilde{A}_{i}$ in Theorem 3.1, respectively, for $i=1,2$.
@@ 4 Proofs
In this section we prove Theorem 3.1. Compared with the exponential case, a major difficulty comes from the fact that $Z^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions. This method is called a BAR approach, and has been used in [1, 2, 11]. Recently, this approach is further studied in [12].
The proof of Theorem 3.1 is composed of three parts. We first introduce Palm distributions and the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
@@ 4.1 Palm distribution and BAR
where $1_{A}$ is the indicator function of the event $A\in{\mathcal{F}}$.
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [12]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S\equiv\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
@@
where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$, and define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
@@
We further define difference operators $\Delta_{e}$ and $\Delta_{s}$ as
@@
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{s}$. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
@@ 4.2 Preliminary lemmas
We will prove Theorem 3.1 using the moment generating functions $\varphi_{i}(\theta_{2})$ and $\varphi_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$. Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
$\varphi_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
$\varphi_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
@@
We note the following lemma, which is obtained from Lemma 5.8 of [2].
@@
Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L(^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L(^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
@@
$\Delta^{(r)}_{i}=\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b\not=0,\end{cases}$ (4.68)

### v4

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation. It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [2, 3, 10, 11]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
@@
11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan.
@@ 1 Introduction
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. It may be considered to approximate using a quasi-birth-and-death process, and it is not hard to numerically compute its performance. However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential. Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@ 2 Framework for the 2-level GI/G/1 queues
We take the stochastic framework introduced in [11] to study the 2-level $GI/G/1$ queue. Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining workload of a customer being served, respectively, at time $t$. Note that the remaining service time cannot be used for $R_{s}(t)$ here because the service rate may change in future. Define
@@ 3.1 Assumptions
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov. Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with respect to the time-shift operator $\Psi$. Let $b_{1}\in\mathbb{R}$ and $b_{2},\ell_{0}>0$ be constants, and assume the following assumptions.
@@
Furthermore, define $\rho_{i}=\lambda_{i}/(c_{i}\mu)$, then (3.b) and (3.c) imply that $1-\rho^{(r)}_{i}=rb_{i}/c_{i}+o(r)$, $\rho_{i}=1$ and $\lambda_{i}=c_{i}\mu$. Hence, $\lambda_{1}=\lambda_{2}$ if and only if $c_{1}=c_{2}$.
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
@@ 3.2 Main results
    $\lim_{r\downarrow 0}\nu_{i|cd}^{(r)}(B)=\widetilde{\nu}_{i|cd}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.6)
@@
$\rho^{(r)}=\alpha^{(r)}_{e}\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}S^{(r)}(k).$ (3.12)
@@
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.13)
@@
Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{\exp}$ which is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
and $\widetilde{A}_{\exp,2}=1-\widetilde{A}_{\exp,1}$.
@@
Since $\sigma_{i}^{2}=2$ and $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$ in the exponential case, $\widetilde{\nu}_{\exp}$ and $\widetilde{A}_{\exp,i}$ agree with $\widetilde{\nu}$ and $\widetilde{A}_{\exp,i}$ in Theorem 3.1, respectively, for $i=1,2$.
@@ 4 Proofs
In this section we prove Theorem 3.1. Compared with the exponential case, a major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [2, 3, 10]. Recently, this approach is further studied in [11].
The proof of Theorem 3.1 is composed of three parts. We first introduce the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
@@ 4.1 Palm distribution and BAR
where $1_{A}$ is the indicator function of the event $A\in{\mathcal{F}}$. For $v=e,s$, $\mathbb{P}^{(r)}_{v}$ is called a Palm distribution concerning the counting process $N^{(r)}_{v}$.
The Palm distributions play a major role for the derivation of the stationary equations in our analysis (e.g., see (4.1)). Intuitively, $\mathbb{P}^{(r)}_{v}$ is considered as the conditional probability distribution given the event that $N^{(r)}_{v}$ counts arrivals for $v=e$ (service completion for $v=s$) at time 00. This event occurs with probability 00 under the stationary framework, so such an event is often handled as a boundary condition in the stationary equation (e.g, see [4]). On the other hand, Palm distributions are defined through time averages under the stationary framework. This enables us to derive the stationary equation without any boundary condition, where the boundary conditions are handled by Palm distributions in the stationary equation. This is quite convenient because the boundary conditions may be only asymptotically satisfied. This is exactly our case.
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [11]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
@@
where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$.
In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit. Define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
@@
For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
@@
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
@@ 4.2 Preliminary lemmas
We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{2})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$. Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
@@
Note that $\eta^{(r)}_{i}(\theta)$ is the value of the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ at $e^{-\theta}$, so it is infinitely many differentiable function of $\theta$. Obviously, $\zeta^{(r)}(\theta)$ has the same property. Hence, twice differentiating the equations in (4.8), we can compute the Taylor expansions of them around the origin, which yields the following lemma. It is formally obtained in [3, Lemma 5.8].
@@
Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
@@
$\Delta^{(r)}_{i}=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
@@
$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (4.68)